a group of 65 people attend a ball game. there were four times as many children as adults in the group. write a system of equations that you could use to set up this problem, where a is the number of adults and c is the number of children in the group. solve the system of equations for c.
Answer by vasu2qute(17)

solution:
No: of adults = a
No: of children =c
children = 4 times adults
c = 4xa=4a
total=65
a+c=65
a+4a=65
5a=65
a=65/5=13
c=4a=4x13=52